Question 2 Tries remaining: Suppose the price was $34 per jacket. Calculate whether there is a surplus or shortage of jackets at that price and Points out of 7.70 Suppose the demand for jackets was given by: QD- 190 -0.2P. The supply of jackets is given by: Qs- 2P -30. the quantity of jackets associated with the surplus or shortage. Enter your response as Qs- Negative numbers indicate shortage. Positive numbers indicate surplus. (Round to the nearest two decimal places if necessary.) Flag question Answer: Check
Answer #1

Answer= - 145.2

Since the demand for jacket is Qd= 190-0.2 Pand supply of jackets is Qs = 2P-30. Equilibrium will occur at the intersection of both the curve.

Therefore, Qd = Qs

190- 0.2P = 2P - 30

220/ 2.2 = P

Equilibrium price will be $100

The Demand at the price $34 is 183.2

The supply at the price $34 is 38

So in the market there exists shortage. Demand exceeds supply.

Qs - Qd= - 145.2